# On the mechanics of inhaled bronchial transmission of pathogenic microdroplets generated from the upper respiratory tract, with implications for infection onset

### v1

@@ III.3 Heaviside trend governing bronchial deposition and penetration
Collapsing the net particulate transmission efficiencies from Table 1, labeled as $\eta$ (in %), on to a log-log scale with particulate diameters $d$ (in $\mathrm{\mu m}$) placed along the horizontal axis, reveals (as is expected) a high lower airway penetration for smaller particulate sizes, followed by a gently sloping dip; see Fig. 5. The behavior could be approximately mathematized with a modified Heaviside function [40] of the following analytical form:
$\eta(d)=\mathcal{C}_{1}+e^{\mathcal{C}_{2}/\left[1+e^{2k\left(d+\mathcal{C}_{3}\right)}\right]},$ (17)
where $k$ and $\mathcal{C}_{i}$, with $i\in\{1,2,3\}$, are constant fitting parameters. For the solid black curve plotted in Fig. 5 and fitting through the data-points from the numerical experiments, the following values were used:
$\mathcal{C}_{1}=-1.95,\quad\mathcal{C}_{2}=4.95,\quad\mathcal{C}_{3}=-25.0,\quad k=0.08.$ (18)
Figure 6: (a) Simulated vorticity contour map on a representative two-dimensional cross-section (the S-ROAM domain) running approximately midway through the cavity; see Fig. 3b. The region with the dominant vortex core is marked by $\mathbb{V}$. (b-h) show two sample particle pathlines (in red) against the S-ROAM streamlines (in grey), respectively for particle sizes 5, 10, 15, 20, 25, 30, and 50 $\mathrm{\mu m}$. The black lines on either side of the S-ROAM domain mark the slip wall boundaries. In each case, the vortex patch $\mathbb{V}$ is mimicked by a set of five point vortices embedded on a the two-dimensional flow field with background unidirectional flow in the $\beta$-direction. The representative particle entering near the middle of the inlet face is marked as $\mathbb{I}$ and the particle entering near the right edge of the model channel is marked as $\mathbb{II}$.
@@ III.4 Analytical projections for particulate transport: consistent with the full-scale numerical findings
Figure 6a demonstrates the vorticity field mapped over the section shown in Fig. 3c. In the reduced-order analytical setup, the vortex patch marked by $\mathbb{V}$ is modeled with a spatial assembly of five point vortices, as described in §II.2. In panels b-h of Fig. 6, the grey curves show the streamlines of the background flow field in the S-ROAM, with the red curves tracing the sample particulate trajectories. Smaller particulates, owing to their low inertia and smaller $St$, follow the streamlines (on which they were embedded at entry points) more closely. For larger particulates, the inertia-dominated dynamics would inhibit downwind penetration through biasing their motion toward wall deposition, e.g., see panels g and (especially) h.
Note that the particulates entering through the middle of the inlet face of the S-ROAM are marked as $\mathbb{I}$ while those entering near the right edge are marked as $\mathbb{II}$. As shown in Fig. 7, the particulates $\mathbb{II}$, being nearer to the vortex region, are deviated more (from the airflow streamlines they were embedded on at the channel inlet) compared to the particulates $\mathbb{I}$. The deviation generically grows as the particulate diameters are increased, which aligns with the findings in Table 1. A greater deviation implies that the particulates are being increasingly shifted toward the channel walls, which in turn would result in a declining deposition and penetration levels in the lower airway.
Figure 7: Comparison of the absolute deviation ($\tilde{\mathbb{D}}$, normalized with respect to the S-ROAM channel width) of the particle pathlines from the respective streamline they were embedded on at the inlet face. The normalized deviation is measured along the $\alpha$-direction (see Fig. 6). Deviation curves for particles entering near the middle of the S-ROAM inlet face is marked as $\mathbb{I}$ and for the particles entering near the right edge of the model channel is marked as $\mathbb{II}$.
IV Open questions: Perspectives on enhancing the biophysical realism of the modeling approach
@@ IV.1 On the interfacial mechanics at mucociliary layers during inhalation
While the numerical and analytical models presented here conclude that aerosols/droplets of the length scale $\mathcal{O}$($10^{1}$) $\mathrm{\mu m}$, formed through shearing and disintegration of intra-URT mucosal filaments, can have significant deposition and penetration in the deep lungs (e.g., see Table 1); to enhance the physiological realism of the model—it is essential that we obtain the actual size distribution and formation rate of liquid particulates generated through viscoelastic separation as the inhaled air brushes past the URT mucus. The present study employed the conservative estimate that 1 particulate of each tested size is formed during each breathing cycle and is a limitation. The reader should however note that unlike expulsion regimes (during exhalation), where imaging-based data collection is relatively straightforward with human subjects expelling particulates into the outside air for different speech and breathing parameters, the current problem of characterizing the internal reverse transport (inhalation) into the lower airway could be somewhat challenging.
Figure 8: (a) Nasopharynx (compare with labels on Fig. 1), which is the initial trigger site for SARS infection owing to the propensity of ciliated epithelial cells with angiotensin-converting enzyme 2 (ACE2) surface receptors that the spike protein viruses bind to during cell intrusion. (b) Representative planar outline of the surface topology at the nasopharynx. Phase 1 stands in for the inhaled air flux, phase 2 signifies the mucus film. The blue arrow is a cartooned depiction of the inhaled airflow profile sweeping over the mucosal topography (in pale red). The black curve in panel (b) delineating the mucosal surface topography is extracted from CT data.
With in vivo tests in human subjects all but unlikely to receive institutional review board approvals, the approach could be essentially two-fold with synergistic numerical modeling and experimental visualizations. One can consider a 2-phase interaction on anatomically realistic upper airway surfaces; the two phases being the mimicked versions of inhaled air (phase 1) and the relatively static mucosal film (phase 2); see Fig. 8. The interfacial mechanics leading to particulate formation and release, along with their spatiotemporal growth and size distribution, could then be assessed with multiphase simulation techniques. While similar problems in a different context have been studied before—the three-dimensional wavy anatomic topology of the URT surface can potentially reveal interesting nuances related to the dynamics. Experimental visualizations with air flux brushing past a viscoelastic film mimicking mucus rheology should then be able to validate the micro-scale mechanics revealed by the computational approach.
@@ IV.2 On experimental validation of deep lung penetration through in vitro physical tests
To verify the numerical modeling of the overall spatial transport, experiments could be conducted in 3D-printed anatomically realistic airway casts that are internally coated with (say) concentrated glycerol (with tuned dilution levels) standing in for mucus. Controlled air flux, with embedded aerosols, could be passed through the cast via suction mechanism set up with a vacuum pump. One approach (amongst others) for getting highly-resolved data on the lower airway penetration could be to use the sophisticated gamma scintigraphy technique [41], wherein the aerosolizing solution to be administered into the three-dimensional cast (with the incoming air) would be seeded with a mildly radioactive element (e.g., Technetium). After the particulates have landed along the airway walls, the radioactive signals (emitted by the deposited mass and hence presenting a measure of local penetration) could be compared with the in silico deposition patterns along the bronchial pathways. Note that for this approach to be functionally consistent, the printed cast should be separable into (at least) two parts along the sagittal mid-section to facilitate controlled liquid coating of its internal surfaces. The separable 3D-printed pieces will also logistically facilitate radio-signal imaging of the wall deposits.
V Conclusion: the main inferences
@@
The particulate transmission trends (see Table 1) from the numerical experiments performed in this study are found to be consistent with the first-principles reduced-order analytical findings that modeled the impact of dominant intra-airway vortex instabilities in the laryngotracheal domain on local particle transport (Fig. 6). Clearly, the results establish that not only the aerosols (i.e., particulates with diameters $\lesssim$ 5 $\mathrm{\mu m}$) but also droplets as large as 10 and 15 $\mathrm{\mu m}$ can exhibit remarkable efficiency at reaching the bronchial spaces and deep lungs, so long as they are sheared away from the URT surface and are still airborne as they enter the larynx. This stands in contrast to the general perception that only the particulates that are smaller than 5 $\mathrm{\mu m}$ can penetrate to the lower airway. Missing the nuance therein is the fact that the incumbent perspective is based on the mechanics of particulates inhaled from the outside air, and our understanding of intra-URT mucosal break-up during inhalation and the subsequent aerial advection profiles in the downwind tract is still nascent.
Figure 9: (a) Sagittal cutting plane through the anatomical airway to extract the intersected curve shown in panel (b) The latter is used to evaluate the tortuosities between P-R and Q-R tracts of the airway.
@@
For a physics-based rationale to explain the derived deposition and penetration profiles, let us consider the tortuosity $\mathcal{T}$ of the pathways to be traversed by a particulate inhaled from outside (called hereafter $P_{o}$), compared to a particulate that is generated from the URT and is still airborne in the laryngeal airspace (hereafter called $P_{u}$). Figure 9 illustrates a representative vertical cutting plane used to extract the tortuosity measurements. For $P_{o}$, the corresponding tortuosity is the ratio of the curved path length and the linear distance in space between the points P and R; let us represent it as $\mathcal{T}_{P_{o}}$. For $P_{u}$, it is similarly the ratio of the curved path length and the linear distance in space between the points Q and R; let us represent it as $\mathcal{T}_{P_{u}}$. The geometric measurements of the spatial distances return: $\mathcal{T}_{P_{o}}\approx 1.90$ and $\mathcal{T}_{P_{u}}\approx 1.04$. Consequently, if the mechanics of $P_{o}$ is inertia-dominated (true for particulates $\gtrsim$ 5 $\mathrm{\mu m}$), they exhibit less success at navigating the highly tortuous pathway (which would have otherwise led them to the lower airway) and are deposited along the anterior URT. On the contrary, particulates of similar sizes, if they only have to traverse the Q-R tract (see Fig. 7), the less tortuous pathway ensures that a high percentage of them would end up depositing along (or, penetrating to) the bronchial domains. As an aside, also note that the tortuosity estimates obtained here match exactly with our previously published data on mammalian airway morphology with different test airway cavities [42]; see panel (d) in Fig. 1 of the cited article. Thus, the findings from this study satisfactorily address the question $Q_{1}$ posed in the introduction (§I), a variant of which is the question heading this subsection.
@@ V.2 How do the transmitted viral loads compare to the infectious dose of the test pathogen?
Infectious dose, $I_{D}$, of a virus quantifies the minimum number of virions that can potentially launch infection in an exposed subject [43, 29] and is a fundamental virological parameter. Independent studies by us [8, 44] and others [9, 45] have verified that $I_{D}\approx$ 300, for SARS-CoV-2 onset. Thus, evidently (per Table 2), the viral load transmitted by, e.g., the 15-$\mathrm{\mu m}$ droplets would alone exceed the $I_{D}$ threshold, thereby providing a mechanics-based rationale for the fast disease progression to the lower airway. The brisk pace is otherwise difficult to explain based exclusively on tissue level proliferation and direct deep lung inhalation of sub 5-$\mathrm{\mu m}$ particulates from outside. However, from a translational paradigm, it is critical to note that the present in silico framework does not yet take into account the host innate and adaptive immune responses to the invading virions. The immunological considerations, once incorporated into the mechanics paradigm, can help rationalize the varying rates of clinical prognosis recorded in different subjects [38]; e.g., the deep lung infection for SARS-CoV-2 has historically ensued over a range of $\mathrm{2-8}$ days in different subjects following the appearance of initial symptoms along the upper respiratory tract. In summary, the findings do satisfactorily address the question $Q_{2}$ posed in §I, a variant of which is the question heading this subsection.

### v2

@@ III.3 Heaviside trend governing bronchial deposition and penetration
Collapsing the transmission efficiencies from Table 1 and Fig. 5a on a 2D plane with particulate diameters $d$ (in $\mathrm{\mu m}$) along the horizontal axis reveals an S-trend with (as expected) a high bronchial penetration for smaller particulate sizes, followed by a gently sloped dip; see Fig 5c. The behavior could be approximated with a modified Heaviside function [61] of the following empirical form:
$\eta(d)=\mathcal{C}_{1}+e^{\mathcal{C}_{2}/\left[1+e^{k\left(d+\mathcal{C}_{3}\right)}\right]},$ (17)
where $k$ and $\mathcal{C}_{i}$, with $i\in\{1,2,3\}$, are constant fitting parameters. The adopted multivariate Heaviside formalism is particularly suitable for modeling the inverted S-shaped trend of bronchial deposition and penetration as a function of the particulate sizes, owing to its ability to represent sharp yet smooth transitions between high and low deposition rates across different size thresholds. This function effectively captures the nonlinear behavior implicit to respiratory systems, where smaller particulates penetrate deeply into the bronchial pathways, while larger ones tend to deposit more proximally or are filtered out. The mathematical ansatz allows for adjustable steepness and transition points, rendering it flexible enough to fit the specific inflection points associated with particle size changes. Herein, the solid red curve, fitting through the numerical data-points, was derived using the Nelder-Mead simplex algorithm for systematic error minimization [62, 63]. The algorithm, also known as the downhill simplex method, is an iterative, heuristic optimization technique designed for unconstrained optimization problems, especially where derivative information is not available or the function is noisy or discontinuous. The curve in Fig 5c invokes the following fitting parameter values:
$\mathcal{C}_{1}=-2.1706,\quad\mathcal{C}_{2}=4.6755,\quad\mathcal{C}_{3}=-23.9592,\quad k=0.23468.$ (18)
@@ III.4 Analytical projections for particulate transport: consistent with the full-scale numerical findings
Fig 6a demonstrates the vorticity field mapped over the section shown in Fig 4c. In the reduced-order analytical setup, the vortex patch core marked by $\mathbb{V}$ is modeled with a spatial assembly of five point vortices, as described in the methods. In Fig 6b-h, the grey curves show the streamlines of the background flow field in the S-ROAM, with the red curves tracing the sample particulate trajectories. Smaller particulates, owing to their low inertia and smaller $St$, follow the streamlines (on which they were embedded at entry points) more closely. For larger particulates, inertia-dominated dynamics, combined with gravitational impaction, tend to inhibit downwind penetration by biasing their motion toward deposition on the walls (see panels g and especially h, in Fig 6). However, some sufficiently large particles, such as those with diameters of 10 and 15 $\mathrm{\mu m}$ (see Fig 6c-d), can still effectively maneuver around the vortex trap to reach the lower airspace. This trend aligns with the full-scale numerical projections.
Note that the particulates entering through the middle of the inlet face of the S-ROAM are marked as $\mathbb{I}$ while those entering near the right edge are marked as $\mathbb{II}$. As shown in Fig 7, the particulates $\mathbb{II}$, being nearer to the vortex region, are deviated more (from the airflow streamlines they were embedded on at the channel inlet) compared to the particulates $\mathbb{I}$. The deviation generically grows as the particulate diameters are increased, which aligns with the numerical findings reported in Table 1. A greater deviation implies that the particulates are being increasingly shifted toward the channel walls, which in turn would result in declining deposition and penetration levels in the downwind lower airway.
Figure 6: Simulation-guided analytical modeling in the laryngotracheal space. (a) Simulated vorticity contour map on a representative two-dimensional cross-section (the S-ROAM domain) running approximately midway through the cavity; see Fig 4b. The core of the dominant vortex patch is conservatively bounded in the rectangle, marked by $\mathbb{V}$, with its area equating the vortex patch area $A$ used in the S-ROAM. Panels (b-h) show two sample particle pathlines (in red) against the S-ROAM streamlines (in grey), respectively for particle sizes 5, 10, 15, 20, 25, 30, and 50 $\mathrm{\mu m}$. The black lines on either side of the S-ROAM domain mark the slip wall boundaries. In each case, the vortex patch $\mathbb{V}$ is mimicked by a set of five point vortices embedded on the two-dimensional flow field with background unidirectional flow in the $\beta$-direction. The representative particles entering near the middle of the inlet face and near the right edge of the model channel are respectively marked as $\mathbb{I}$ and $\mathbb{II}$.
Figure 7: Particle transport trend in the analytical case. Comparison of the absolute deviation ($\tilde{\mathbb{D}}$, normalized with respect to the S-ROAM channel width) of the particle pathlines from the respective streamline they were embedded on at the inlet face. The deviation is measured along the $\alpha$-direction (see Fig 6). Deviation curves for particles entering near the middle and the right edge of the model channel inlet are respectively marked as $\mathbb{I}$ and $\mathbb{II}$.
IV Discussion: perspectives on enhancing the biophysical realism of the modeling approach
@@ IV.1 On the interfacial mechanics at mucociliary layers during inhalation
While the numerical and analytical models presented here conclude that particulates of the length scale $\mathcal{O}$($10^{1}$) $\mathrm{\mu m}$, formed through shearing and disintegration of intra-URT mucosal filaments, can indeed have significant deposition and penetration along the bronchial tubes and in the deeper lung recesses (see Table 1); to enhance the physiological realism of the model—it is essential that we obtain the actual size distribution and formation rate of the liquid particulates generated through viscoelastic separation as the inhaled air brushes past the URT mucus. The present study used the conservative estimate that 1 particulate of each size tested is formed during each breathing cycle and is a limitation.
The reader should however note that unlike expulsion regimes (during exhalation), where imaging-based data collection is relatively straightforward with human subjects expelling particulates into the outside air for different speech and breathing parameters, the current problem of characterizing the internal reverse transport (inhalation) into the lower airway in live subjects could be somewhat challenging. The approach could consequently be two-fold with synergistic numerical modeling and experimental visualizations. One can consider a 2-phase interaction on anatomically realistic upper airway surfaces; the two phases being the mimicked versions of inhaled air (phase 1) and the relatively static mucosal substrate (phase 2); see Fig 8a, and explore the interfacial mechanics leading to particulate formation and release, along with their generation rate, spatiotemporal growth, and size distribution.
@@ IV.2 On experimental validation of deep lung penetration through in vitro physical tests
To verify the numerical modeling of the overall spatial transport, further experiments could be conducted within 3D-printed anatomically realistic airway casts internally coated with (say) concentrated glycerol (with tuned dilution levels) standing in for mucus. Controlled air flux, with embedded aerosols, could be passed through the cast via suction mechanism set up with a vacuum pump. One approach (among others) to obtain highly resolved data on lower airway penetration could be to use the sophisticated gamma scintigraphy technique [64], wherein the solution to be aerosolized and administered into the cast (with the incoming air) would be seeded with a mildly radioactive element (e.g., Technetium). After the particulates have landed along the airway walls, the radioactive signals (emitted by the deposited mass and thus presenting a measure of local penetration) could be compared with the in silico deposition patterns along the bronchial pathways. In this context, the reader may scan our previous works with experimental validations [64, 65]; in particular, [64] called on gamma scintigraphy measurements within anatomically accurate and transparent 3D-printed airway casts to physically verify numerically modeled sinonasal deposition from over-the-counter nasal sprays. The modeling protocol therein implemented similar computational schemes as in the present analysis.
IV.3 On limitations implicit to the modeling framework and its clinical relevance
In terms of achieving true biological realism, it is important to note the limitations implicit to the numerical modeling and theoretical analysis described in this study. The approach here algorithmizes the complex dynamics of infection onset, particularly omitting the role of immune responses and mucosal properties. In an actual infection setting, the innate and adaptive immune responses would be activated, likely leading to dynamic alterations in mucus properties that could impact intra-URT particulate formation, clearance, and the pathogen’s state within these fragments. By focusing on a static view of mucosal fragmentation and particulate generation, this study currently does not account for such dynamic physiological responses that may influence the outcomes of viral transport and downwind deposition.
Figure 8: Interfacial interactions and morphological complexity along mucus-coated upper airway walls. (a) Representative planar outline of the nasopharyngeal surface topology. Phase 1 = inhaled air; Phase 2 = mucus substrate. The blue arrow indicates the inhaled air flux sweeping over the mucus (in pale red). (b) Tortuosity measurement of the anatomical space using a sagittal cutting plane.
Next, the reported results, while establishing the basic plausibility of the URT-to-LRT transmission mechanism through the use of a test case, are for a specific anatomical geometry and baseline parameters that do not (yet) explicitly account for patient-to-patient variability. Realistic differences in airway anatomy, mucosal characteristics, and immune response between individuals are factors that could lead to different fragmentation and transport dynamics. Future work should incorporate sensitivity studies that test the robustness of the modeling approach across a range of physiological conditions and individual variability.
Moreover, the “so what” question remains. The findings, while precise in a computational sense and confirming the possibility of rapid lung invasion orchestrated by URT-derived pathogen-laden particulates, do not yet translate into actionable clinical recommendations. Future research should assess whether the insights gained could inform clinical practices or support diagnostic tools, possibly through personalized respiratory physics models that account for specific physiological traits. Such models could potentially identify subjects at higher risk of clinical morbidity, e.g., from brisk onset of severe lung infection.
V Conclusion: the main takeaways
@@
The bronchial transmission trends from the numerical experiments (see Table 1 and Fig 5) are found to be consistent with the first-principles reduced-order analytical findings that modeled the impact of dominant intra-airway vortex instabilities in the laryngotracheal domain on local particle transport (Fig 6). Not only the aerosols (i.e., particulates with diameters $\lesssim$ 5 $\mathrm{\mu m}$) but also droplets as large as 10 and 15 $\mathrm{\mu m}$ exhibit remarkable efficiency at reaching the bronchial spaces and deep lungs, so long as they are sheared away from the URT surface and are still airborne as they enter the larynx. This contradicts the general perception that only particulates smaller than 5 $\mathrm{\mu m}$ can comfortably penetrate the lower airway. Missing the nuance therein is the fact that the incumbent perspective is based on the mechanics of particulates inhaled from the outside air, and our understanding of intra-URT mucosal breakup during inhalation and the subsequent aerial advection in the downwind tract is still nascent.
@@
For a physics-based rationale to explain the derived deposition and penetration profiles, let us consider the tortuosity $\mathcal{T}$ of the pathways to be traversed by a particulate inhaled from outside (called hereafter $P_{o}$), compared to a particulate that is generated within the URT and is still airborne in the laryngeal airspace (hereafter called $P_{u}$). Fig 8b illustrates a representative vertical cutting plane used to extract the tortuosity measurements. For $P_{o}$, the corresponding tortuosity is the ratio of the curved path length and the linear distance in space between the points P and R; let us represent it as $\mathcal{T}_{P_{o}}$. For $P_{u}$, it is similarly the ratio of the curved path length and the linear distance in space between the points Q and R; let us represent it as $\mathcal{T}_{P_{u}}$. The geometric measurements return: $\mathcal{T}_{P_{o}}\approx 1.90$ and $\mathcal{T}_{P_{u}}\approx 1.04$. Consequently, if the mechanics of $P_{o}$ is inertia-dominated (true for particulates $\gtrsim$ 5 $\mathrm{\mu m}$), they exhibit less success at navigating the highly tortuous pathway and are deposited along the anterior URT; see Fig 3 in this context. On the contrary, particulates of similar sizes, if they only have to traverse the Q-R tract, the less tortuous pathway ensures that a high percentage of them would end up reaching the bronchial domains; see Fig 5a-b. As an aside, also note that the tortuosity estimates obtained here match exactly with our previously published data on mammalian airway morphology [66]; see panel (d) in figure 1 of the cited paper. Thus, the findings of this study satisfactorily address the question $Q_{1}$ (see introduction).
@@ V.2 How do the transmitted viral loads compare to the infectious dose of the test pathogen?
Infectious dose, $I_{D}$, of a virus quantifies the minimum number of virions that can potentially launch infection in an exposed subject [67, 49] and is a fundamental virological parameter. Independent studies by us [25, 26] and others [24, 68] have verified that $I_{D}\approx$ 300, for SARS-CoV-2. Thus, evidently (per Table 2), the viral load transmitted by, e.g., the 15-$\mathrm{\mu m}$ droplets would alone exceed the $I_{D}$ threshold, thereby providing a mechanics-based rationale for the fast disease progression to the lower airway. The brisk pace is otherwise difficult to explain based exclusively on tissue level proliferation and direct deep lung inhalation of dominantly sub 5-$\mathrm{\mu m}$ particulates from outside. However, from a translational perspective, it is critical to note that the present in silico framework does not yet take into account the host innate and adaptive immune responses to the invading virions which can lead to shifts in mucus properties altering breakup, while affecting the dynamics of clearance and state of the virus in the URT-derived fragments. The immunological considerations, once incorporated into the mechanics paradigm, can help rationalize the varying rates of clinical prognosis recorded in different subjects [59]; e.g., deep lung infection for SARS-CoV-2 has historically ensued over a range of $\mathrm{2-8}$ days following the appearance of initial URT symptoms. In summary, the findings do satisfactorily address the question $Q_{2}$ posed in the introduction.